# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 6.4.2. The divisor $\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}$
for $i\in\mathbb{Z}/N\mathbb{Z}$, and that $\mathcal{D}_{\mathcal{X},i,i+1}\rightarrow S_{\mathcal{X}}$ is the Mumford toric degeneration defined by the polyhedral decomposition of $C([v_{i},v_{i+1}])$ given in Equation (6.7), and the PL function $\varphi_{i}$. In particular, the general fiber of $\mathcal{D}_{\mathcal{X},i,i+1}\rightarrow S_{\mathcal{X}}$ is $\mathbb{P}^{1}$, and its central fiber over $0_{\mathcal{X}}$ is a chain of $\mathbb{P}^{1}$-components indexed by the integral points $v\in[v_{i},v_{i+1}]\cap P_{\mathbb{Z}}$. Hence the result follows. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Then, the divisors $\mathcal{D}_{\mathcal{X}^{\prime}}$ in the mirror families $\mathcal{Y}_{\mathcal{X}^{\prime}}\rightarrow S_{\mathcal{X}^{\prime}}$ corresponding to the projective crepant resolutions $\mathcal{X}^{\prime}\rightarrow\overline{\mathcal{X}}$ glue into a divisor $\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}$ in the glued mirror family $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}\rightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ over the movable secondary fan.
@@
This is immediate by the gluing of theta functions in Equation (6.4). ∎
@@ 6.5. Torus action and extension over the bogus cones
Let $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}})\rightarrow S_{\mathcal{X}}^{\mathrm{msec}}$ be the polarized family as in Equation (6.5), $\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}}$ the theta divisor as in Definition 6.21, and $\mathcal{D}_{\mathcal{X}}^{\mathrm{sec}}$ the divisor as in Lemma 6.25. In this section, we describe how to canonically extend the family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})$ defined over $S_{\mathcal{X}}^{\mathrm{msec}}$, to a family over the toric stack with stacky fan given by the stacky “secondary fan” defined as follows.
First, recall that the “bogus cones” of the Mori fan $MF(\mathcal{X}/\overline{\mathcal{X}})$, as discussed in §6.4, are of the form $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})+\langle\mathrm{Ex}(f^{\prime})\rangle$, for a birational map $f^{\prime}:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$, where $\mathcal{X}^{\prime}$ is $\mathbb{Q}$-factorial. We call two bogus cones $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})+\langle\mathrm{Ex}(f^{\prime})\rangle$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})+\langle\mathrm{Ex}(f^{\prime\prime})\rangle$ equivalent if the smallest cone of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$ containing $\mathrm{Nef}(\mathcal{X}^{\prime})$ equals the smallest cone of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$ containing $\mathrm{Nef}(\mathcal{X}^{\prime\prime})$.
@@
[…]
@@
For every bogus cone $\sigma=\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})+\langle\mathrm{Ex}(f^{\prime})\rangle$, let $M_{\sigma}$ be the sublattice of $\mathrm{Span}(\sigma)\cap\mathrm{Pic}(\mathcal{X}/\overline{\mathcal{X}})$ spanned by $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})\cap\mathrm{Pic}(\mathcal{X}/\overline{\mathcal{X}})$ and by the classes of irreducible components of $\mathcal{X}_{0}$ contracted by $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$. Equivalent bogus cones define the same sublattices, and so we have a well-defined sublattice $M_{\sigma}$ for every cone $\sigma$ of the secondary fan. The data of the sublattices $M_{\sigma}$ is a toric stacky data as in [94, Definition 4.1], and so defines a stacky fan $\mathcal{S}ec(\mathcal{X}/\overline{\mathcal{X}})$ with underlying fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$.
We denote from now on by $S^{\mathrm{sec}}_{\mathcal{X}}$ the proper toric variety whose fan is the secondary fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$, and by $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ the proper toric Deligne–Mumford stack with stacky fan $\mathcal{S}ec(\mathcal{X}/\overline{\mathcal{X}})$ – see [94, §4.1] and [38] for the general theory of stacky fans and toric stacks. In what follows, we show that the family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})$ extends to a family $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})$ over $\mathcal{S}^{sec}_{\mathcal{X}}$. To do this, we first describe a natural torus action on $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})$. Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Denote by $T_{0}$ the torus $\mathrm{Spec}\,\mathbb{C}[\mathbb{Z}^{P_{\mathbb{Z}}}]$. Note that $\mathbb{Z}^{P_{\mathbb{Z}}}$ can be naturally identified with the group of divisors of $\mathcal{X}$ supported on $\mathcal{X}_{0}$, since the points of $P_{\mathbb{Z}}$ are in a natural one-to-one correspondence with the irreducible components of $\mathcal{X}_{0}$. In particular, for every divisor $F$ supported on $\mathcal{X}_{0}$, we have a corresponding one-parameter subgroup $\mathbb{C}^{\star}\subset T_{0}$. There is a natural action of $T_{0}$ on $S_{\mathcal{X}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$ such that, for every divisor $F$ supported on $\mathcal{X}_{0}$, the corresponding one-parameter subgroup acts with weights $F\cdot\beta$ on the monomials $t^{\beta}$.
Lemma 6.27.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. There exists a unique action of the torus $T_{0}$ on the mirror family
$\mathcal{Y}_{\mathcal{X}}=\mathrm{Proj}\,\mathcal{R}_{\mathcal{X}}\longrightarrow S_{\mathcal{X}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]\,,$

### v3

@@ 6.4.4. The divisor $\mathcal{D}_{\mathcal{X}}$
for $i\in\mathbb{Z}/N\mathbb{Z}$, and that $\mathcal{D}_{\mathcal{X},i,i+1}\rightarrow S_{\mathcal{X}}$ is the Mumford toric degeneration defined by the polyhedral decomposition of $C([v_{i},v_{i+1}])$ given in Equation (6.9), and the piecewise linear function $\varphi_{i}$. In particular, the general fiber of $\mathcal{D}_{\mathcal{X},i,i+1}\rightarrow S_{\mathcal{X}}$ is $\mathbb{P}^{1}$, and its central fiber over $0_{\mathcal{X}}$ is a chain of $\mathbb{P}^{1}$-components indexed by the integral points $v\in[v_{i},v_{i+1}]\cap P_{\mathbb{Z}}$. Hence the result follows. ∎
Theorem 6.36.
Proof.
First note that $(\mathcal{Y}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}})\rightarrow S_{\mathcal{X}}$ is flat since the algebra $\mathcal{R}_{\mathcal{X}}$ of theta functions is a free $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$-module. On the other hand, by Theorem 6.30, the algebra $\mathcal{R}_{\mathcal{X}}$ is finitely generated, and so the morphism $\mathcal{Y}_{\mathcal{X}}=\mathrm{Proj}\,\mathcal{R}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ is projective.
To show that the fiber $(Y_{t},D_{t})$ over $t\in S_{\mathcal{X}}$ is semi-log-canonical and $K_{Y_{t}}+D_{t}=0$, consider the smallest toric stratum $\mathrm{Spec}\,\mathbb{C}[H]$ of $S_{\mathcal{X}}$ containing $t$, corresponding to a face $H$ of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. In particular, $t$ is contained in the dense torus orbit $T_{H}:=\mathrm{Spec}\,\mathbb{C}[H^{\mathrm{gp}}]$ in $\mathrm{Spec}\,\mathbb{C}[H]$. As in the proof of Lemma 6.29, there is a corresponding birational contraction $f_{H}:\mathcal{X}\rightarrow\mathcal{X}_{H}$ such that $H=NE(\mathcal{X}/\mathcal{X}_{H})$, and the restriction $\mathcal{Y}_{\mathcal{X}}|_{T_{H}}$ is independent of the choice of the crepant resolution $\mathcal{X}\rightarrow\mathcal{X}_{H}$. Hence, by Lemma 6.20, we can assume that there exists a good divisor $F$ on $\mathcal{X}$ over $\mathcal{X}_{H}$. Let $G$ be the face of $H$ spanned by the curves $C$ with $F\cdot C=0$, and denote $J:=H\setminus G$. Consider the one-parameter subgroup $\mathbb{C}^{\star}\subset T_{0}$ associated with $F$ as in §6.6. Consider the corresponding orbit $\mathbb{C}^{\star}\subset T_{H}$ passing through $t$. Its closure in $\mathrm{Spec}\,\mathbb{C}[H]$ is an affine line $\mathbb{A}^{1}$ with $0\in\mathrm{Spec}\,\mathbb{C}[G]\simeq\mathrm{Spec}\,\mathbb{C}[H]/J$. By Lemma 6.26, the restriction of $(\mathcal{Y}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}})$ to this line is a $\mathbb{C}^{\star}$-equivariant family. Hence, by the proof of [46, Lemma 8.33], it suffices to show Theorem 6.36 for the central fiber of this family over $0\in\mathrm{Spec}\,\mathbb{C}[G]\simeq\mathrm{Spec}\,\mathbb{C}[H]/J$. By Lemma 6.28, this central fiber is a fiber of a Mumford toric degeneration. Hence, the result follows from [5, Theorem 1.2.14]. ∎
6.5. Polarized mirror over the movable secondary fan
Let $\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration. We constructed in §6.4 a mirror family $(\mathcal{Y}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}}+\epsilon\mathcal{C}_{\mathcal{X}})\rightarrow S_{\mathcal{X}}$, where $0<\epsilon<\!\!<1$. In this section, we glue together the mirror families corresponding to the various crepant resolutions of $\mathcal{X}^{\mathrm{can}}$.
By Proposition 3.36, the contraction $\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is a relative Mori dream space. Hence, the corresponding Mori fan $MF(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ is a complete fan in $\mathrm{Pic}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})_{\mathbb{R}}$, whose maximal dimensional cones are indexed by the finitely many $\mathbb{Q}$-factorial birational models $f^{\prime}:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ of $\mathcal{X}$ over $\mathcal{X}^{\mathrm{can}}$, and are given by $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})+\langle\mathrm{Ex}(f^{\prime})\rangle$, where $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$ is the relative nef cone of $\mathcal{X}^{\prime}$ over $\mathcal{X}^{\mathrm{can}}$, and $\langle\mathrm{Ex}(f^{\prime})\rangle$ is the non-negative span of the $f^{\prime}$-exceptional divisors. If $\langle\mathrm{Ex}(f^{\prime})\rangle\neq\{0\}$, that is, if at least one divisor is contracted, the corresponding cone in $MF(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ is referred to as bogus. In other words, a cone is non-bogus if $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ is small, that is, by Theorem 6.7, if $\mathcal{X}^{\prime}\rightarrow\mathcal{X}^{\mathrm{can}}$ is a projective resolution, and $\mathcal{X}^{\prime}\rightarrow\Delta$ is an open Kulikov degeneration. The union of non-bogus cones is the movable cone $\mathrm{Mov}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$.
Following [53, §2.2], we define the movable secondary fan $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ as a coarsening of the restriction $\mathrm{MovF}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ of the Mori fan to the movable cone $\mathrm{Mov}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. Two maximal cones $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\mathcal{X}^{\mathrm{can}})$ of $\mathrm{MovF}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, related by a flop $b:\mathcal{X}^{\prime}\dashrightarrow\mathcal{X}^{\prime\prime}$ are called equivalent if the corresponding piecewise linear functions $\varphi^{\prime\prime}-\nu_{b}\circ\varphi^{\prime}$ in §6.2 is equal to zero.
@@ 6.5. Polarized mirror over the movable secondary fan
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. There exists a unique fan $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ with support the movable cone $\mathrm{Mov}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, whose maximal cones are the unions of equivalent maximal cones of $\mathrm{MovF}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$.
@@
This follows from the proof of [53, Theorem 2.12]. ∎
Lemma 6.38.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Two maximal cones $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\mathcal{X}^{\mathrm{can}})$ of $\mathrm{Mov}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ are equivalent if and only if $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ are connected by a sequence of multiple M1 flops.
Proof.
If $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ are connected by a sequence of multiples M1 flops, then the maximal cones $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\mathcal{X}^{\mathrm{can}})$ are equivalent by Lemma 6.12.
Conversely, if $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\mathcal{X}^{\mathrm{can}})$ are equivalent, then, by Lemma 6.37, $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\mathcal{X}^{\mathrm{can}})$ are both contained in a maximal cone $C$ of $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. As $C$ is convex, there exists a path contained in $C$, starting in $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$, ending $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\mathcal{X}^{\mathrm{can}})$, and only intersecting cones of $MF(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ of codimension zero or one. Therefore, it suffices to prove the result when $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\mathcal{X}^{\mathrm{can}})$ share a common codimension one face, that is, when $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ are related by a flop of relative Picard rank one, which by Theorem 6.7 is either a multiple M1 or a multiple M2 flop. By Lemma 6.13, $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ would not be equivalent if they were related by a multiple M2 flop, and so we conclude that $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ are related by a multiple M1 flop. ∎
Theorem 6.39.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration. Then, the mirror families $(\mathcal{Y}_{\mathcal{X}^{\prime}},\mathcal{D}_{\mathcal{X}^{\prime}}+\epsilon\,\mathcal{C}_{\mathcal{X}^{\prime}})\rightarrow S_{\mathcal{X}^{\prime}}$ and the relatively ample line bundles $\mathcal{L}_{\mathcal{X}^{\prime}}$ corresponding to the projective crepant resolutions $\mathcal{X}^{\prime}\rightarrow\mathcal{X}^{\mathrm{can}}$ uniquely glue into a family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}_{\mathcal{X}}^{\mathrm{msec}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})\rightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ and a relatively ample line bundle $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$ over the toric variety $S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ with fan given by the movable secondary fan $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$.
Proof.
Let $C$ be a maximal cone of $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. By Lemma 6.38 and Theorem 6.14, all mirror algebras $\mathcal{R}_{\mathcal{X}^{\prime}}$ with $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})\subset C$ have the same structure constants $N_{pqr}^{\beta}(\mathcal{X}^{\prime})$. In particular, these structure constants are only non-zero if $\beta\in NE(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$ for every $\mathcal{X}^{\prime}$ with $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})\subset C$, that is if $\beta\in C^{\vee}_{\mathbb{Z}}$, where $C^{\vee}$ is the dual cone of $C$. Hence, the mirror families $(\mathcal{Y}_{\mathcal{X}^{\prime}},\mathcal{D}_{\mathcal{X}^{\prime}}+\epsilon\,\mathcal{C}_{\mathcal{X}^{\prime}})\rightarrow S_{\mathcal{X}^{\prime}}$ and line bundles $\mathcal{L}_{\mathcal{X}^{\prime}}$ glue into a mirror family $(\mathcal{Y}_{C},\mathcal{D}_{C}+\epsilon\,\mathcal{C}_{C})\rightarrow\mathrm{Spec}\,\mathbb{C}[C^{\vee}_{\mathbb{Z}}]$ and a line bundle $\mathcal{L}_{C}$, where $\mathcal{Y}_{C}:=\mathrm{Proj}\,\mathcal{R}_{C}$ and
If $C$ and $C^{\prime}$ are two maximal cones of $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, with corresponding sections $\varphi_{C}$ and $\varphi_{C^{\prime}}$ as in (6.4), and related by a flop $b$, then it follows from Theorem 6.14 that the map
$\vartheta_{p}$ $\longmapsto t^{\nu_{b}(\varphi_{C}(p))-\varphi_{C^{\prime}}(p)}\vartheta_{p}\,$,
is an isomorphism of $\mathbb{C}[(C\cap C^{\prime})^{\vee}_{\mathbb{Z}}]$-algebras. If $C$, $C^{\prime}$, and $C^{\prime\prime}$ are three maximal cones of $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, then the cocycle condition $\psi_{C,C^{\prime\prime}}=\psi_{C,C^{\prime}}\circ\psi_{C^{\prime},C^{\prime\prime}}$ holds. Therefore, using the isomorphisms $\psi_{C,C^{\prime}}$, one can glue the polarized families $(\mathcal{Y}_{C},\mathcal{D}_{C},\mathcal{L}_{C})\rightarrow\mathrm{Spec}\,\mathbb{C}[C^{\vee}_{\mathbb{Z}}]$ into a polarized family
$\nu_{\mathrm{msec}}:(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}_{\mathcal{X}}^{\mathrm{msec}}\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}})\longrightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$
over the toric variety $S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ defined by the movable secondary fan $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. By Equation (6.12), for every $p\in(hP)_{\mathbb{Z}}$, $\vartheta_{p}$ naturally becomes a section of the line bundle $(\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}})^{\otimes h}\otimes\nu_{\mathrm{msec}}^{\star}L_{p}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}$, where $L_{p}$ is the line bundle on $S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ defined by the 1-cocycle $t^{\nu_{b}(\varphi_{C}(p))-\varphi_{C^{\prime}}(p)}$. By Lemma 6.13, we have always $\nu_{b}(\varphi_{C}(p))-\varphi_{C^{\prime}}(p)=0$ for $p\in P_{\mathbb{Z}}$, and so the line bundle $L_{p}$ is trivial for every $p\in P_{\mathbb{Z}}$. In particular, all the theta functions $\vartheta_{p}$ with $p\in P_{\mathbb{Z}}$ are sections of $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$, and so the theta divisors $\mathcal{C}_{C}$ glue into a theta divisor
$\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}}:=\left\{\sum_{p\in P_{\mathbb{Z}}}\vartheta_{p}=0\right\}\in|\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}|\,.$
@@ 6.6. Torus action and extension over the bogus cones
In this section, we describe how to canonically extend the family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})\rightarrow S_{\mathcal{X}}^{\mathrm{msec}}$ given by Theorem 6.39, to a family over the toric stack with stacky fan given by the stacky “secondary fan” defined below following [53].
First, recall that the “bogus cones” of the Mori fan $MF(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, as discussed at the beginning of §6.5, are of the form $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})+\langle\mathrm{Ex}(f^{\prime})\rangle$, for a birational map $f^{\prime}:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$, where $\mathcal{X}^{\prime}$ is $\mathbb{Q}$-factorial. We call two bogus cones $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})+\langle\mathrm{Ex}(f^{\prime})\rangle$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\mathcal{X}^{\mathrm{can}})+\langle\mathrm{Ex}(f^{\prime\prime})\rangle$ equivalent if the smallest cone of $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ containing $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$ equals the smallest cone of $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ containing $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\mathcal{X}^{\mathrm{can}})$.
@@
The secondary fan of $\mathcal{X}/\mathcal{X}^{\mathrm{can}}$, denoted by $\mathrm{Sec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ is the complete fan in $\mathrm{Pic}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})_{\mathbb{R}}$, obtained by adding to the movable secondary fan $\mathrm{MovSec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ the maximal cones given by the unions of equivalent bogus cones of the Mori fan $MF(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$.
@@
For every bogus cone $\sigma=\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})+\langle\mathrm{Ex}(f^{\prime})\rangle$, let $M_{\sigma}$ be the sublattice of $\mathrm{Span}(\sigma)\cap\mathrm{Pic}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ spanned by $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})\cap\mathrm{Pic}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ and by the classes of irreducible components of $\mathcal{X}_{0}$ contracted by $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$. Equivalent bogus cones define the same sublattices, and so there exists a well-defined sublattice $M_{\sigma}$ for every cone $\sigma$ of the secondary fan. The data of the sublattices $M_{\sigma}$ is a toric stacky data as in [98, Definition 4.1], and so defines a stacky fan $\mathcal{S}ec(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ with underlying fan $\mathrm{Sec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. We denote from now on by $S^{\mathrm{sec}}_{\mathcal{X}}$ the proper toric variety whose fan is the secondary fan $\mathrm{Sec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, and by $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$ the proper toric Deligne–Mumford stack with stacky fan $\mathcal{S}ec(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ – see [98, §4.1] and [41] for the general theory of stacky fans and toric stacks. In what follows, we show that the family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})$ extends to a family $(\mathcal{Y}^{\mathrm{sec}}_{\mathcal{X}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})$ over $\mathcal{S}^{sec}_{\mathcal{X}}$.
The action of the torus $T_{0}$ given by Lemma 6.26 on the affine toric varieties $S_{\mathcal{X}^{\prime}}$ corresponding to the projective crepant resolutions $\mathcal{X}^{\prime}\rightarrow\mathcal{X}^{\mathrm{can}}$ naturally induces an action of $T_{0}$ on the toric varieties $S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ and $S^{\mathrm{sec}}_{\mathcal{X}}$, and on the toric stack $\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$. It follows from the Equation (6.12) that the actions of $T_{0}$ on the families $\mathcal{Y}_{\mathcal{X}^{\prime}}\to S_{\mathcal{X}^{\prime}}$ given by Lemma 6.26 glue into an action of $T_{0}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}\rightarrow S_{\mathcal{X}}^{\mathrm{msec}}$.
Theorem 6.41.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration. There exists a unique $T_{0}$ equivariant extension of the family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})\rightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ to a family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}^{\mathrm{sec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}^{\mathrm{sec}}_{\mathcal{X}})\to\mathcal{S}^{\mathrm{sec}}_{\mathcal{X}}$.